### v1

@@ 2.2 Tropical tape recorder
At low oxygenation states (<1<1% PAL), seasonal differences in temperature and the Brewer-Dobson circulation are still present, but their magnitude is smaller so the tropical tape recorder effect is muted. Whilst the H2O mixing ratios still vary, the tape recorder effect is no longer visible from the scale shown in Fig. 5 at 0.1% PAL because there is no clear periodic signal. We verified that the PI simulation has a tape recorder and the 0.1% PAL simulation does not have one using a Fourier analysis. Therefore, the seasonal mechanism which transports more water vapour into the stratosphere during warmer periods in the TTL is no longer effective in the 0.1% PAL case.
@@ 2.3 Total hydrogen mixing ratio
Figure 7: Top: $f_{\textrm{T}}(H)$, the total hydrogen mixing ratio is given in parts per million by volume (ppmv) and is shown against pressure for the 150% PAL (grey), PI (black), 50% PAL (dark green), 10% PAL (green), 5% PAL (light green), 1% PAL (dark blue), 0.5% PAL (blue), and 0.1% PAL (light blue) simulations. $f_{\textrm{T}}(H)$ is calculated using Eq. (2). Hence, the mixing ratio contribution from H, H2O, H2, and CH4 is multiplied by a factor of 1, 2, 2, and 4, respectively. Bottom left: The total hydrogen mixing ratio contribution from H (dash dotted), H2 (dashed), H2O (unbroken), and CH4 (dotted) is shown for the PI and 10% PAL simulations and bottom right: the 1% PAL and 0.1% PAL simulations. In Eq. (2), the mixing ratios of H2, H2O, and CH4, are multiplied by 2, 2, and 4, respectively, to determine their contribution to $f_{\textrm{T}}(H)$.
@@ 3.1 The oxygen-ozone valve
Our work shows that O2 levels during the Proterozoic may have partially controlled the diffusion-limited flux for hydrogen escape. However, this will depend on several other factors, such as the continental distribution and the Earth’s obliquity which can affect atmospheric dynamics and seasonal cycles, respectively. The fainter Sun, the albedo of the Earth through time, and various greenhouse gases (including CH4) which can act to cool the stratosphere but warm the surface (2024PNAS..12119228L), will have also impacted TTL temperatures and hydrogen escape rates. For instance, a cooling of the stratosphere can increase the amount of O3, just like the current increasing atmospheric content of CO2 is aiding a possible, albeit complex, recovery of the O3 layer (2002JGRD..107.4049R; fahey2018scientific; 2022WCD.....3..139I; 2021EGUGA..23.5376M).
@@
There exist estimates of many of these aforementioned variables through Earth’s history, but each of them come with uncertainties and vary throughout the Proterozoic. liu2025evolution suggest that the O3 layer could have been kept at low levels due to protracted atmospheric iodine concentrations throughout the Proterozoic. Atmospheric iodine would destroy O3 through catalytic cycles (WACCM6 does not currently include iodine in its chemical mechanism). The computed WACCM6 O3 columns are lower than the baseline estimates from liu2025evolution, but when liu2025evolution included substantial atmospheric iodine (e.g. $20\times$ modern iodine concentrations), then their O3 columns are lower than the WACCM6 results. If we incorporated iodine, then this could lead to a reduced rate of predicted hydrogen escape at a given O2 concentration due to diminished O3 heating.
@@ 3.2 Future work for Earth
Our atmospheric predictions are relevant for both Earth’s past and future climate states. However, this statement comes with several caveats. The moist-adiabatic lapse rate (the rate at which the temperature of a parcel of saturated air decreases with altitude as it rises in the atmosphere under adiabatic conditions), strongly depends on temperature. Thus, the total hydrogen mixing ratio in the troposphere of the Earth is determined by the incoming radiation, albedo, and greenhouse effect. For modern Earth, these parameters are relatively well known. We did not attempt to fully account for all climatic changes since the Proterozoic began; instead, as a first step, we focused on the effect on the escape rate by changing O2 only. The range of O2 concentrations we accounted for are relevant to the past 2.4 billion years and are not relevant to the Archean eon (4–2.4 Gyr ago) when molecular oxygen was much lower than present day and Proterozoic (2.4–0.541 Gyr ago) concentrations (2020SciA....6.1420C). The different climate states that might have existed throughout the Proterozoic and Phanerozoic [^5] By climate states, here we refer to the variety of factors that would have affected weather, atmospheric circulation systems, and surface temperature. These include glaciations and hotter periods, as well as the various continental land masses, fainter Sun, and greenhouse gas concentrations. (0.541 Gyr ago–present day) should be simulated in future work. A hotter troposphere may have existed due to increased amounts of CH4 or CO2, or both. In this case, the warmer TTL would have increased hydrogen escape rates by elevating the amount of H2O reaching the stratosphere. On the other hand, colder temperatures during Earth’s glacial periods would cool the tropopause (2019JGRD..12411819G) and reduce the stratospheric H2O concentration.
@@
Several calculations of the lifespan of the Earth’s biosphere have been attempted (1982Natur.296..561L; 1992Natur.360..721C; 2000TellB..52...94F; von2003biogenic; 2018AsBio..18..469R; 2021NatGe..14..138O; 2023IJAsB..22..272M; 2024arXiv240910714G). For instance, 2021NatGe..14..138O calculated that in $1.08\pm 0.14$ billion years ($1\ \sigma$), Earth’s atmospheric O2 will drop to 1% PAL. More recently, 2024arXiv240910714G suggested that Earth’s biosphere may last for up to 1.6–1.86 Gyr. Most of these estimates for biospheric termination are before the Earth is predicted to lose its water and become, by the most basic definition, uninhabitable. As O2 levels drop following the extinction of photosynthesizing life (2021NatGe..14..138O), O3 will decrease, and the upper troposphere and TTL may be colder than under otherwise equivalent conditions, diminishing the amount of hydrogen available in the upper atmosphere which can escape to space and cause irreversible water loss (the future rates of hydrogen escape would likely have a negligible effect on oxidizing the Earth to reverse the loss of O2 from ceasing photosynthesis). Hence, our results suggest that the ocean loss timescale could be extended. Thus, investigations aiming to assess the lifetime of Earth’s future habitability should incorporate a 3D chemistry-climate model that includes comprehensive atmospheric oxygen chemistry and that can adequately represent the 3D movement of air parcels through the TTL. Additionally, it is possible that our WACCM6 simulations could be used as input to a model which has a higher top (lower minimum pressure), such as the WACCM-X model (2018JAMES..10..381L), or even used as an input into a computational model that explicitly calculates atmospheric escape (e.g., 2018PNAS..115..260D; 2023MNRAS.523..286S).
@@ 3.3 Future work for exoplanets
There are various planetary system parameters which affect the climate state of exoplanets. Crucial to the atmospheric dynamics is the planetary rotation rate and whether it is tidally locked (2014RSPTA.37230084F; 2016A&A...596A.112T; 2017A&A...601A.120B; 2019AsBio..19...99D; 2025PSJ.....6....5B). The circulation affects the atmospheric temperature structure and how tracers such as H2O are advected. Other salient factors include the strength and spectral shape of the incoming radiation (2020A&A...639A..99E), the continental distribution or lack thereof (2018ApJ...854..171L; 2021ApJ...910L...8Z; 2022MNRAS.513.2761M), obliquity (2019ApJ...877L...6K), eccentricity (2023MNRAS.524.1491L), flares (2021NatAs...5..298C; 2023MNRAS.518.2472R), simulation resolution (2020ApJ...894...84S; 2021ApJ...913..101L; 2022ApJ...934..149S; 2022ApJ...940...87K; 2023NatAs...7.1070Y; 2024arXiv240219277S; 2024ApJ...965....5G), as well as the planetary Bond albedo and surface composition (2011ApJ...729...54C; 2019ApJ...884...75D; 2020MNRAS.495....1M). Ocean salinity can also be important in modulating the climate of rocky planets (2022GeoRL..4995748O).
@@
[…]
[Conclusions] Since the dawn of the Proterozoic (2.4 Gyr ago), several properties of the Earth have changed. The continental configuration has altered several times, the Sun’s luminosity has increased, and the concentration of atmospheric gases such as O2, CO2, and CH4 have fluctuated.
@@
Between 0.1% PAL and 150% PAL of O2, the upward diffusion of H2O is modulated in a non-linear manner, with 5% and 10% PAL having the maximum predicted diffusion-limited escape rate of hydrogen due to the peak in O3 concentrations causing localized heating. The predicted escape rate changes by up to a factor of 4.7 between all of the simulations, meaning that whilst the oxygen valve affects diffusion-limited hydrogen escape, in the absence of other factors, it would not have resulted in relatively high levels of hydrogen escape and thus water loss since the GOE.
@@
The time-averaged data is currently available in the Dryad data repository. The data with time-variability will be made available upon acceptance of the manuscript.
@@ Acknowledgements.
We would like to thank the reviewers for the first round of reviews which helped to improve the manuscript. G.J.C. thanks the Science and Technology Facilities Council for financial support during the PhD when the simulations were conducted (grant number ST/T506230/1). G.J.C. thanks James Rogers and Oli Shorttle for helpful discussions regarding Earth’s history and hydrogen escape. F. Sainsbury-Martinez and C. Walsh would like to thank UK Research and Innovation for support under grant number MR/T040726/1. Additionally, C. Walsh would like to thank the University of Leeds and the Science and Technology Facilities Council for financial support (ST/X001016/1). M.B. appreciates support from a CSH Postdoctoral Fellowship. This work was undertaken on ARC4, part of the High Performance Computing facilities at the University of Leeds, UK. The CESM project is supported primarily by the National Science Foundation (NSF). This material is based upon work supported by the National Center for Atmospheric Research (NCAR), which is a major facility sponsored by the NSF under Cooperative Agreement 1852977.

### v2

@@ 2.2 Tropical tape recorder
At low oxygenation states (<1<1% PAL), seasonal differences in temperature and the Brewer-Dobson circulation are still present, but their magnitude is smaller so the tropical tape recorder effect is muted. Whilst the H2O mixing ratios still vary, the tape recorder effect is no longer visible from the scale shown in Fig. 5 at 0.1% PAL because there is no clear periodic signal. Variance-preserving Fourier spectra show that the annual harmonic (1 cycle yr-1) dominates the stratospheric H2O variability in the PI simulation between 100 and 30 hPa. In contrast, the 0.1% PAL simulation exhibits a much weaker annual contribution relative to total variance at these pressure levels, consistent with the absence of a tape-recorder signal. Therefore, the seasonal mechanism which transports more water vapour into the stratosphere during warmer periods in the TTL is no longer effective in the 0.1% PAL case.
@@ 2.3 Total hydrogen mixing ratio
Figure 7: Top: $f_{\textrm{T}}(H)$, the total hydrogen mixing ratio is given in parts per million by volume (ppmv) and is shown against pressure for the 150% PAL (grey), PI (black), 50% PAL (dark green), 10% PAL (green), 5% PAL (light green), 1% PAL (dark blue), 0.5% PAL (blue), and 0.1% PAL (light blue) simulations. $f_{\textrm{T}}(H)$ is calculated using Eq. (2). Hence, the mixing ratio contribution from H, H2O, H2, and CH4 is multiplied by a factor of 1, 2, 2, and 4, respectively. Bottom left: The total hydrogen mixing ratio contribution from H (dash dotted), H2 (dashed), H2O (unbroken), and CH4 (dotted) is shown for the PI and 10% PAL simulations and bottom right: the 1% PAL and 0.1% PAL simulations.
@@ 3.1 The oxygen-ozone valve
Our work shows that O2 levels during the Proterozoic may have partially controlled the diffusion-limited flux for hydrogen escape. However, this will depend on several other factors, such as the continental distribution and the Earth’s obliquity which can affect atmospheric dynamics and seasonal cycles, respectively. The fainter Sun, the albedo of the Earth through time, and various greenhouse gases (including CH4) which can act to cool the stratosphere but warm the surface (2024PNAS..12119228L), will have also impacted TTL temperatures and hydrogen escape rates. For instance, a cooling of the stratosphere can increase the amount of O3, just like the current increasing atmospheric content of CO2 is aiding a possible, albeit complex, recovery of the O3 layer (2002JGRD..107.4049R; fahey2018scientific; 2022WCD.....3..139I; 2021ACP....2111041M).
@@
There exist estimates of many of these aforementioned variables through Earth’s history, but each of them come with uncertainties and vary throughout the Proterozoic. liu2025evolution suggest that the O3 layer could have been kept at low levels due to protracted atmospheric iodine concentrations throughout the Proterozoic. Atmospheric iodine would destroy O3 through catalytic cycles (WACCM6 does not currently include iodine in its chemical mechanism). The computed WACCM6 O3 columns are lower than the baseline estimates from liu2025evolution, but when liu2025evolution included substantial atmospheric iodine (e.g. $20\times$ modern marine iodine concentrations that was emitted to the atmosphere), then their O3 columns are lower than the WACCM6 results. If we incorporated iodine, then this could lead to a reduced rate of predicted hydrogen escape at a given O2 concentration due to diminished O3 heating.
3.2 Model limitations
Our atmospheric predictions are relevant for both Earth’s past and future climate states. However, this statement comes with several caveats. The moist-adiabatic lapse rate (the rate at which the temperature of a parcel of saturated air decreases with altitude as it rises in the atmosphere under adiabatic conditions), strongly depends on temperature. Thus, the total hydrogen mixing ratio in the troposphere of the Earth is determined by the incoming radiation, albedo, and greenhouse effect. For modern Earth, these parameters are relatively well known. We did not attempt to fully account for all climatic changes since the Proterozoic began; instead, as a first step, we focused on the effect on the escape rate by changing O2 only. The range of O2 concentrations we accounted for are relevant to the past 2.4 billion years and are not relevant to the Archean eon (4 – 2.4 Gyr ago) when molecular oxygen was much lower than present day and Proterozoic (2.4 – 0.541 Gyr ago) concentrations (2020SciA....6.1420C). We remind the reader that WACCM6 is designed for the modern Earth and the current set of simulations do not account for a faster rotation rate in the past (1987PreR...37...95Z; 2016GeoRL..43.5716B; laskar2023did) or different continental coverage (2021NRvEE...2..358M). Additionally, we note that changes to the orbital parameters, rotation rate, or obliquity, may affect the tape recorder signal.
Since 2022RSOS....911165C was published, ji2023comparison showed that absorption of light by H2O and CO2 in the S-R bands, as well as scattering in the S-R bands, are increasingly important at lower atmospheric O2 mixing ratios. The simulations presented in this work do not include either of these treatments. Some WACCM6 simulations (at 10% PAL, 1% PAL, and 0.1% PAL) with H2O and CO2 absorption in the S-R bands were included in ji2023comparison, and we also ran some simulations with fixed surface fluxes (e.g. for CH4). Whilst the absolute mixing ratios do deviate when this absorption is included, the effect of changing O2 on temperature is stronger. In terms of scattering, ji2023comparison showed that it is most important at $\leq 1$% PAL of O2. Inclusion of scattering will decrease the total O3 column and likely result in a cooler TTL region. Again, the conclusion that hydrogen escape can be be modified by O2 concentrations remains unaffected, especially when the prescribed CH4 mixing ratio sets a minimum on the hydrogen escape rate because it is not cold-trapped. Nevertheless, future simulations where WACCM6 includes both absorption and scattering in the S-R bands will provide more comprehensive atmospheric calculations.
@@ 3.3 Future work for Earth
The different climate states that might have existed throughout the Proterozoic and Phanerozoic [^5] By climate states, here we refer to the variety of factors that would have affected weather, atmospheric circulation systems, and surface temperature. These include glaciations and hotter periods, as well as the various continental land masses, fainter Sun, and greenhouse gas concentrations. (0.541 Gyr ago – present day) should be simulated in future work. A hotter troposphere may have existed due to increased amounts of CH4 or CO2, or both. In this case, the warmer TTL would have increased hydrogen escape rates by elevating the amount of H2O reaching the stratosphere. On the other hand, colder temperatures during Earth’s glacial periods would cool the tropopause (2019JGRD..12411819G) and reduce the stratospheric H2O concentration.
@@
Several calculations of the lifespan of the Earth’s biosphere have been attempted (1982Natur.296..561L; 1992Natur.360..721C; 2000TellB..52...94F; von2003biogenic; 2018AsBio..18..469R; 2021NatGe..14..138O; 2023IJAsB..22..272M; Graham_2024_Extension). For instance, 2021NatGe..14..138O calculated that in $1.08\pm 0.14$ billion years ($1\ \sigma$), Earth’s atmospheric O2 will drop to 1% PAL. More recently, Graham_2024_Extension suggested that Earth’s biosphere may last for up to 1.6 – 1.86 Gyr. Most of these estimates for biospheric termination are before the Earth is predicted to lose its water and become, by the most basic definition, uninhabitable. As O2 levels drop following the extinction of photosynthesizing life (2021NatGe..14..138O), O3 will decrease, and the upper troposphere and TTL may be colder than under otherwise equivalent conditions, diminishing the amount of hydrogen available in the upper atmosphere which can escape to space and cause irreversible water loss (the future rates of hydrogen escape would likely have a negligible effect on oxidizing the Earth to reverse the loss of O2 from ceasing photosynthesis). Hence, our results suggest that the ocean loss timescale could be extended. Thus, investigations aiming to assess the lifetime of Earth’s future habitability should incorporate a 3D chemistry-climate model that includes comprehensive atmospheric oxygen chemistry and that can adequately represent the 3D movement of air parcels through the TTL. Additionally, it is possible that our WACCM6 simulations could be used as input to a model which has a higher top (lower minimum pressure), such as the WACCM-X model (2018JAMES..10..381L), or even used as an input into a computational model that explicitly calculates atmospheric escape (e.g., 2018PNAS..115..260D; 2023MNRAS.523..286S).
@@ 3.4 Future work for exoplanets
The H2O tropical tape recorder is a stratospheric signal of seasonality, impressing the varying TTL temperatures on the stratospheric H2O mixing ratio. Prior work showed how a 15% reduction in global O3 (Xie_2008_O3_tape) could cool the TTL and raise it in height, ultimately weakening the tape recorder signal. Here we find this effect exacerbated further as O2 and O3 reduce by orders of magnitude. Transmission spectra of exoplanet atmospheres probe the stratosphere, and whilst beyond the detection limit of today’s telescopes, such chemical variability may be detectable with future technology or where exoplanets exhibit higher levels of variation that Earth (e.g., for exoplanets with eccentricity $\geq 0.4$ 2023MNRAS.524.1491L) or different types of stratospheric variability (Cohen_2022_LASO). Seasonality of biosignatures and other gases (Olson_2018_season), combined with the H2O tape recorder signal, may aid in determining whether a planet is inhabited. Should there be no tape recorder signal due to a tropopause that does not significantly vary with seasons, then this variability will be more difficult to observe, and H2O will be harder to constrain due to a dry middle-upper atmosphere.
@@
There are various planetary system parameters which affect the climate state of exoplanets. Crucial to the atmospheric dynamics is the planetary rotation rate and whether it is tidally locked (2014RSPTA.37230084F; 2016A&A...596A.112T; 2017A&A...601A.120B; 2019AsBio..19...99D; 2025PSJ.....6....5B). The circulation affects the atmospheric temperature structure and how tracers such as H2O are advected. Other salient factors include the strength and spectral shape of the incoming radiation (2020A&A...639A..99E), the continental distribution or lack thereof (2018ApJ...854..171L; 2021ApJ...910L...8Z; 2022MNRAS.513.2761M), obliquity (2019ApJ...877L...6K), eccentricity (2023MNRAS.524.1491L), flares (2021NatAs...5..298C; 2023MNRAS.518.2472R), simulation resolution (2020ApJ...894...84S; 2021ApJ...913..101L; 2022ApJ...934..149S; 2022ApJ...940...87K; 2023NatAs...7.1070Y; Sergeev_2024_Impact; 2024ApJ...965....5G), as well as the planetary Bond albedo and surface composition (2011ApJ...729...54C; 2019ApJ...884...75D; 2020MNRAS.495....1M). Ocean salinity can also be important in modulating the climate of rocky planets (2022GeoRL..4995748O).
@@
4 Conclusions
Since the dawn of the Proterozoic (2.4 Gyr ago), several properties of the Earth have changed. The continental configuration has shifted several times, the Sun’s luminosity has increased, and the concentration of atmospheric gases such as O2, CO2, and CH4 have fluctuated.
@@ 4 Conclusions
Between 0.1% PAL and 150% PAL of O2, the upward diffusion of H2O is modulated in a non-linear manner, with 5% and 10% PAL having the maximum predicted diffusion-limited escape rate of hydrogen due to O3 concentrations causing localized heating. The predicted escape rate changes by up to a factor of 4.7 between all of the simulations, meaning that whilst the oxygen valve affects diffusion-limited hydrogen escape, in the absence of other factors, it would not have resulted in relatively high levels of hydrogen escape and thus water loss since the GOE.
@@
The time-averaged data is currently available in the Dryad data repository.
@@ Acknowledgements.
We would like to thank the reviewers for their comments which helped to improve the manuscript. G.J.C. thanks the Science and Technology Facilities Council for financial support during the PhD when the simulations were conducted (grant number ST/T506230/1). G.J.C. thanks James Rogers and Oli Shorttle for helpful discussions regarding Earth’s history and hydrogen escape. F. Sainsbury-Martinez and C. Walsh would like to thank UK Research and Innovation for support under grant number MR/Z00019X/1 Additionally, C. Walsh would like to thank the University of Leeds and the Science and Technology Facilities Council for financial support (ST/X001016/1). M.B. appreciates support from a CSH Postdoctoral Fellowship. This work was undertaken on ARC4, part of the High Performance Computing facilities at the University of Leeds, UK. The CESM project is supported primarily by the U.S. National Science Foundation.